OpenStudy (anonymous):

A furniture maker is making a chair and needs 12 feet of wood. He has chosen to make the chair out of maple, which costs $25 per foot, and pine, which costs $28 per foot. If the furniture maker wants to spend a total of $25.75 per foot, how many feet of each type of wood will he need? What would be the system of equations I would get from this, so I can solve? I've tried to look at it but can't really get one exactly.

OpenStudy (anonymous):

All I need is help finding the system of equations on here, not the answer so yeah.

OpenStudy (anonymous):

Let's call x feet of maple and y feet of pine, if you need 12 feet of wood x+y=12 On the other hand maples costs 25 $/foot and pine 28 $/foot so if we have 25*x+28*y=25.75 we hace the quantities that makes the final price of 25.75

OpenStudy (anonymous):

have*

OpenStudy (anonymous):

25x+28y=25.75 x+y=12

OpenStudy (kinggeorge):

The last equation should not be \[25x+28y=25.75\]It should rather be \[25x+28y=12(25.75)\]

OpenStudy (anonymous):

True kinggeorge, cause the price is per foot

OpenStudy (anonymous):

Sorry

OpenStudy (anonymous):

Oh okay so then, 25x+28y=309 x+y=12 and no prob. knock, everyone makes mistakes once in a while. I've made mistakes on OS also so yeah.

OpenStudy (anonymous):

Ty @KingGeorge & @knock

OpenStudy (anonymous):

^^ indeed, yw

OpenStudy (kinggeorge):

you're welcome.